OpenStudy (anonymous):

Find an equation in slope-intercept form (where possible) for the line. Through (-4, 2) and (4, -3)

OpenStudy (anonymous):

slop= rise/run = -5/8 y-intercept = -5/8*4+2=-5/2+2 = -1/2 equation = y=-5x/8-1/2